# Active Human Feedback Collection via Neural Contextual Dueling Bandits

### v1

@@ Abstract
Collecting human preference feedback is often expensive, leading recent works to develop principled algorithms to select them more efficiently. However, these works assume that the underlying reward function is linear, an assumption that does not hold in many real-life applications, such as online recommendation and LLM alignment. To address this limitation, we propose Neural-ADB, an algorithm based on the neural contextual dueling bandit framework that provides a principled and practical method for collecting human preference feedback when the underlying latent reward function is non-linear. We theoretically show that when preference feedback follows the Bradley-Terry-Luce model, the worst sub-optimality gap of the policy learned by Neural-ADB decreases at a sub-linear rate as the preference dataset increases. Our experimental results on problem instances derived from synthetic preference datasets further validate the effectiveness of Neural-ADB.
@@ 1 Introduction
Collecting human preference feedback is essential in many real-life applications, like online recommendations (Kohli et al., 2013; Wu et al., 2023; Zhang & Wang, 2023; Yang et al., 2024), content moderation (Avadhanula et al., 2022), medical treatment design (Lai & Robbins, 1985; Bengs et al., 2021), prompt optimization (Lin et al., 2024), and aligning large language models (Bai et al., 2022; Mehta et al., 2023; Menick et al., 2022; Chaudhari et al., 2024; Das et al., 2024; Ji et al., 2024), to ensure systems effectively align with user preferences and exhibit desired behaviors. However, this process is often costly due to the need for skilled evaluators, the complexity of tasks, and the time-intensive nature of producing high-quality, reliable human feedback. To address the challenge of balancing cost and effectiveness in aligning systems, this paper proposes principled and practical algorithms for efficiently collecting human feedback sequentially and adaptively to achieve the desired system behavior. Specifically, this paper aims to answer the following fundamental question:
@@
Recent works (Mehta et al., 2023; Das et al., 2024) have modeled the problem of active human feedback collection as an active version of the contextual dueling bandit problem (ADB for brevity) (Saha, 2021; Bengs et al., 2022; Li et al., 2024), where context-arm pair in the contextual dueling bandits corresponds to a task for which human preference feedback is collected and then proposed algorithms to select context-arm pairs for human feedback sequentially and adaptively by exploiting collected preference dataset, i.e., past context-arm pairs with their preference feedback. The preference feedback between two context-arm pairs is commonly assumed to follow the Bradley-Terry-Luce (BTL) model [^1] For more than two context-arm pairs, preferences are typically modeled using the Plackett-Luce model (Soufiani et al., 2014). (Hunter, 2004; Luce, 2005; Saha, 2021; Bengs et al., 2022; Li et al., 2024; Lin et al., 2024; Verma et al., 2025) in which the probability of preferring a context-arm pair over others is proportional to the exponential of its reward. In many real-life applications, the number of context-arm pairs (e.g., prompt-response pair in prompt optimization or user-movie pair in online movie recommendation) can be large or even infinite. Therefore, the reward for each context-arm pair is assumed to be parameterized by an unknown function of its feature vector.
@@
An efficient arm selection strategy requires estimating the reward function to guide the arm selection process effectively. Since the reward function may not always be linear in practice, this paper parameterizes the reward function via a non-linear function, which needs to be estimated using the available preference dataset. This non-linear function can be estimated by using methods like Gaussian processes (Williams & Rasmussen, 2006; Srinivas et al., 2010) or neural networks (Zhou et al., 2020; Zhang et al., 2021; Verma et al., 2025). However, Gaussian processes have limited expressive power and fail to optimize highly complex functions. In contrast, neural networks (NNs) have greater expressive power, making them well-suited for modeling complex functions (Dai et al., 2023; Lin et al., 2023; 2024; Verma et al., 2025).
In this paper, we propose a neural active contextual dueling bandit algorithm, Neural-ADB, which uses an NN to estimate the unknown reward function using the available preference dataset. The context selection in Neural-ADB is adapted from Das et al. (2024), while arm selection strategies are based on, respectively, upper confidence bound (UCB) and Thompson sampling (TS), and adapted from Verma et al. (2025). Due to the differences in context selection strategy, arm selection strategies, and the use of a non-linear reward function, our theoretical analysis is completely different than related existing work (Mehta et al., 2023; Das et al., 2024). One of the key theoretical contributions of this paper is providing an upper bound on the maximum Mahalanobis norm of a vector from the fixed input space, measured with respect to the inverse of a positive definite Gram matrix that is constructed using finite, adapted samples from that space. Building on this result, we prove that the worst sub-optimality gap (defined in Eq. 1) of the policy learned by Neural-ADB decreases at a sub-linear rate as the preference dataset size increases. Specifically, our contribution are as follows:
  • We introduce the setting of active contextual dueling bandits with a non-linear reward function in Section 2. In Section 3, we propose a practical neural active contextual dueling bandit algorithm, Neural-ADB, which uses an NN to estimate the unknown reward function from the available preference dataset and then uses this estimate into the arm selection strategies.
  • One of the main theoretical contributions of this paper is proving an upper bound on the maximum Mahalanobis norm of a vector from the fixed input space, as measured with respect to the inverse of a positive definite Gram matrix (1), where the gram matrix is constructed using finite, adapted samples from that input space. We show that this upper bound decays at a sub-linear rate as the number of samples used in the Gram matrix increases. This result itself is of independent interest, as it gives valuable insights beyond the specific application of our work.
@@
  • Finally, in Section 4, our experimental results further validate the different performance aspects of Neural-ADB.
@@ 2 Problem Setting
We formulate active human preference feedback collection as an active contextual dueling bandit problem, where a labeler (human or simulator) provides preference feedback for a chosen pair of options (arms), such as prompt-response pairs in prompt optimization (Lin et al., 2024) and LLM alignment (Mehta et al., 2023; Das et al., 2024).
Active contextual dueling bandit problem.   We consider an active contextual dueling bandit problem, where the underlying latent reward function can be non-linear. In each iteration of this problem, the learner’s goal is to select a triplet containing a context and two arms for collecting preference feedback from a labeler/human such that the collected preference dataset leads to superior performance. Let $\mathcal{C}$ be the set of contexts and $\mathcal{A}$ be the set of all possible arms. In each iteration, the learner selects a context $c_{t}\in\mathcal{X}$ and then two arms (denoted as $a_{t,1}$ and $a_{t,2}$) from the set of arms $\mathcal{A}$. After selecting the triplet of context and two arms, the learner receives a stochastic preference feedback $y_{t}$, where $y_{t}=1$ implies the arm $a_{t,1}$ is preferred over arm $a_{t,2}$ for the context $c_{t}$ and $y_{t}=0$ otherwise. We use $\varphi(c_{t},a)$ to denote the context-arm feature vector for context $c_{t}$ and an arm $a$, where $\varphi:\mathcal{C}\times\mathcal{A}\rightarrow\mathbb{R}^{d}$ is a known feature map.
@@
  • Let $\kappa_{\mu}\doteq\inf\limits_{c\in\mathcal{C},a,b\in\mathcal{A}}\dot{\mu}(f(\varphi(c,a))-f(\varphi(c,b)))>0$ for all triplets of context and arms.
@@
Performance measure.  We denote the collected preference dataset up to $T$ iterations by $\mathcal{D}_{T}=\{(c_{s},a_{s,w},a_{s,l},y_{s})\}_{s=1}^{T}$, where $a_{s,w}\succ a_{s,l}$ for the selected context $c_{s}$ in iteration $s$. We aim to learn a policy, $\pi:\mathcal{C}\rightarrow\mathcal{A}$ from the collected preference dataset $\mathcal{D}_{T}$ that achieves the worst suboptimality across all contexts in $\mathcal{C}$, which is defined as follows:
@@ 3 Algorithm for Active Human Preference Feedback Collection
This section introduces Neural-ADB, a simple yet principled and practical algorithm designed to efficiently select context-arm pairs for collecting preference feedback. Neural-ADB consists of two main components: context selection and arm selection. Since the arm selection strategy depends on the estimated reward function, we first explain how an NN can be used to estimate the unknown reward function. We will then give details of the context and arm selection strategies, followed by our theoretical results.
@@ 3.1 Reward function estimation using neural network
where $\text{ReLU}(v)=\max\{v,0\}$, $\bm{W}_{1}\in\mathbb{R}^{w\times d}$, $\bm{W}_{l}\in\mathbb{R}^{w\times w}$ for $2\leq l<D$, $\bm{W}_{D}\in\mathbb{R}^{w\times 1}$. The parameters of the NN are represented by $\theta=\left(\text{vec}\left(\bm{W}_{1}\right);\cdots\text{vec}\left(\bm{W}_{D}\right)\right)$, where $\text{vec}\left(A\right)$ transforms an $m\times n$ matrix $A$ into a vector of dimension $mn$. We use $p$ to represent the total number of NN parameters, which is given by $p=dw+w^{2}(L-1)+w$, and $g(x;\theta)$ to denote the gradient of NN $h(x;\theta)$ with respect to $\theta$. At the end of each iteration $t$, the available preference dataset $\mathcal{D}_{t}=\left\{(c_{s},a_{s,w},a_{s,l},y_{s})\right\}_{s=1}^{t}$ is used to estimate the reward function $f$ by training an NN $h$ (parameterized by $\theta_{t+1}$) using gradient descent to minimize the following binary cross entropy loss function:
@@
where $\theta_{0}$ denotes the initial parameter of the NN that is initialized according to the standard practice in neural bandits (Zhou et al., 2020; Zhang et al., 2021). Minimizing the first term in the above loss function (that involves the summation over the $t$ terms) corresponds to finding the maximum log-likelihood estimate of the parameters $\theta$.
@@ 3.2 Neural-ADB
In this section, we propose a simple yet principled and practical algorithm, Neural-ADB, that consists of two key components: Context selection and arm selection. Neural-ADB works as follows: At the beginning of the iteration $t$, we first select the context by adapting context selection strategy [^3] Note that selecting contexts uniformly at random suffer a constant suboptimality gap (Das et al., 2024, Theorem 3.2). from Das et al. (2024) as follows:
@@
where $V_{t-1}=\frac{\lambda}{\kappa_{\mu}}\mathbb{I}_{p}+\sum_{s=1}^{t-1}z_{s}{z_{s}}^{\top}\frac{1}{w}$ in which $z_{s}=\varphi^{\prime}(c_{s},a_{s,w})-\varphi^{\prime}(c_{s},a_{s,l})=g(\varphi(c_{s},a_{s,w});\theta_{0})-g(\varphi(c_{s},a_{s,l});\theta_{0})$, and $g(\varphi(c_{s},a_{s,i});\theta_{0})/\sqrt{w}$ is used as the Random features approximation for the context-arm feature vector $\varphi(c_{s},a_{s,i})$. After selecting context $c_{t}$, Neural-ADB uses the trained NN (as an estimate for the unknown reward function) to decide which two arms must be selected. To do so, Neural-ADB uses UCB- and TS-based arm selection strategies, which efficiently balance the exploration-exploitation trade-off (Lattimore & Szepesvári, 2020) due to the bandit nature of preference feedback, as it is only observed for the selected arms.
@@
The second arm $a_{t,2}$ is selected optimistically to focus on exploration, i.e., by maximizing the UCB value as follows:
@@
$\sigma_{t-1}^{2}(c,a,b)\doteq\frac{\lambda}{\kappa_{\mu}}\left\|\frac{1}{\sqrt{w}}(\varphi^{\prime}(c,a)-\varphi^{\prime}(c,b))\right\|^{2}_{V_{t-1}^{-1}}.$ (6)
A larger value of $\sigma_{t-1}^{2}(c_{t},a_{t,1},b)$ implies that arm $b$ is significantly different from $a_{t,1}$, given the contexts and arm pairs already selected. As a result, the second term in Eq. 5 makes the second arm different from the first arm.
TS-based arm selection strategy.   Thompson sampling (TS) selects an arm based on its probability of being the best (Agrawal & Goyal, 2013; Li et al., 2024). Several works (Li et al., 2024; Chowdhury & Gopalan, 2017; Agrawal & Goyal, 2013; Chapelle & Li, 2011) have shown that TS empirically outperforms UCB-based bandit algorithms. Therefore, we propose a TS-based arm selection strategy in which the first arm is also selected using Eq. 4 and only the second arm is selected differently. To select the second arm, it first samples a score $s_{t}(b)\sim\mathcal{N}\big{(}h(\varphi(c_{t},b);\theta_{t})-h(\varphi(c_{t},a_{t,1});\theta_{t}),\nu_{T}^{2}\sigma_{t-1}^{2}(c_{t},a_{t,1},b)\big{)}$ for every arm $b\in\mathcal{A}\setminus\{a_{t,1}\}$ and then selects the second arm as follows:
@@ 3.3 Theoretical Results
The proof of the first part follows directly from Theorem 1 of (Verma et al., 2025) while the second follows from Lemma 10 with Eq. (27) of (Verma et al., 2025). Equipped with 1, 1, and 2, we will now provide an upper bound on the worse sub-optimality gap for a policy learned by Neural-ADB while using UCB- and TS-based arm selection strategy for a given context.
@@
The proof follows by applying 2, setting $z^{\prime}=\varphi^{\prime}(c,a^{\star})-\varphi^{\prime}(c,\pi(c))$ in Eq. 6, and then using 1. Note that $\varepsilon^{\prime}_{w,T}=O(1/T)$ and $\widetilde{d}=\widetilde{o}(\sqrt{T})$ as long as the NN width $w$ is large enough (Verma et al., 2025). Above 2 and 3 show that the worst sub-optimality gap of the policy learned by Neural-ADB with UCB- and TS-based arm selection strategies decreases at a sub-linear rate with respect to the size of preference dataset, specifically at rate of $\tilde{O}((\tilde{d}/T)^{\frac{1}{2}})$, where $\tilde{O}$ hides the logarithmic factors and constants. Further, the decay rate of the worst sub-optimality gap for Neural-ADB improves by a factor of $\tilde{O}(({\tilde{d}\log T})^{\frac{1}{2}})$ compared to exiting algorithms (Mehta et al., 2023; Das et al., 2024), thereby bridging the gap between theory and practice.

### v2

# hyperrefIgnoring empty anchor

@@ Abstract
Collecting human preference feedback is often expensive, leading recent works to develop principled algorithms to select them more efficiently. However, these works assume that the underlying reward function is linear, an assumption that does not hold in many real-life applications, such as online recommendation and LLM alignment. To address this limitation, we propose Neural-ADB, an algorithm based on the neural contextual dueling bandit framework that provides a principled and practical method for collecting human preference feedback when the underlying latent reward function is non-linear. We theoretically show that when preference feedback follows the Bradley-Terry-Luce model, the worst sub-optimality gap of the policy learned by Neural-ADB decreases at a sub-linear rate as the preference dataset increases. Our experimental results on preference datasets further corroborate the effectiveness of Neural-ADB.
@@ 1 Introduction
Collecting human preference feedback is essential in many real-life applications, like online recommendations (Kohli et al., 2013; Wu et al., 2023; Zhang and Wang, 2023; Yang et al., 2024), content moderation (Avadhanula et al., 2022), medical treatment design (Lai and Robbins, 1985; Bengs et al., 2021), prompt optimization (Lin et al., 2024), and aligning large language models (Bai et al., 2022; Menick et al., 2022; Mehta et al., 2023; Chaudhari et al., 2024; Das et al., 2024; Ji et al., 2024), to ensure systems effectively align with user preferences and exhibit desired behaviors. However, this process is often costly due to the need for skilled evaluators, the complexity of tasks, and the time-intensive nature of producing high-quality, reliable human feedback. To address the challenge of balancing cost and effectiveness in aligning systems, this paper proposes principled and practical algorithms for efficiently collecting human feedback sequentially and adaptively to achieve the desired system behavior. Specifically, we aim to answer the following fundamental question:
@@
Recent works (Mehta et al., 2023; Das et al., 2024) have modeled the problem of active human feedback collection as an active version of the contextual dueling bandit problem (ADB for brevity) (Saha, 2021; Bengs et al., 2022; Li et al., 2024), where context-arm pair in the contextual dueling bandits corresponds to a task for which human preference feedback is collected and then proposed algorithms to select context-arm pairs for human feedback sequentially and adaptively by exploiting collected preference dataset, i.e., past context-arm pairs with their preference feedback. The preference feedback between two context-arm pairs is commonly assumed to follow the Bradley-Terry-Luce (BTL) model [^1] For more than two context-arm pairs, preferences are typically modeled using the Plackett-Luce model (Soufiani et al., 2014). (Hunter, 2004; Bengs et al., 2022; Li et al., 2024; Lin et al., 2024; Verma et al., 2025) in which the probability of preferring a context-arm pair over others is proportional to the exponential of its reward. In many real-life applications, the number of context-arm pairs (e.g., user-movie pair in online movie recommendation) can be large or even infinite. Therefore, the reward for each context-arm pair is assumed to be an unknown function of its feature vector, such as a linear function (Mehta et al., 2023; Das et al., 2024).
@@
An efficient arm selection strategy requires estimating the reward function to guide the arm selection process effectively. Since the reward function may not always be linear in practice, this paper parameterizes the reward function via a non-linear function, which needs to be estimated from the available preference dataset by using methods like Gaussian processes (Williams and Rasmussen, 2006; Srinivas et al., 2010) or neural networks (Zhou et al., 2020; Zhang et al., 2021). However, Gaussian processes have limited expressive power and fail to optimize highly complex functions. In contrast, neural networks (NNs) have greater expressive power, making them well-suited for modeling complex functions (Dai et al., 2023; Lin et al., 2023, 2024; Verma et al., 2025).
In this paper, we propose a neural active contextual dueling bandit algorithm, Neural-ADB, which uses an NN to estimate the unknown reward function using the available preference dataset. The context selection in Neural-ADB is adapted from Das et al. (2024), while arm selection strategies are based on, respectively, upper confidence bound (UCB) and Thompson sampling (TS), and adapted from Verma et al. (2025). Due to the differences in context selection strategy, arm selection strategies, and the use of a non-linear reward function, our theoretical analysis is completely different than related existing work (Mehta et al., 2023; Das et al., 2024). One of the key theoretical contributions of this paper is providing an upper bound on the maximum Mahalanobis norm of a vector from the fixed input space, measured with respect to the inverse of a positive definite Gram matrix that is constructed using finite, adapted samples from that space. Building on this result, we prove that the worst sub-optimality gap (defined in Eq. 1) of the policy learned by Neural-ADB decreases at a sub-linear rate as the preference dataset size increases.
Specifically, our key contributions can be summarized as follows:
  • We introduce the setting of active contextual dueling bandits with a non-linear reward function in Section 2. In Section 3, we propose a neural active contextual dueling bandit algorithm, Neural-ADB, which uses an NN to estimate the unknown reward function from the available preference dataset and then uses this estimate into the arm selection strategies.
  • We prove an upper bound on the maximum Mahalanobis norm of a vector from the fixed input space, as measured with respect to the inverse of a positive definite Gram matrix (1), where the gram matrix is constructed using finite, adapted samples from that input space. We show that this upper bound decays at a sub-linear rate as the number of samples used in the Gram matrix increases. This theoretical result itself is of independent interest, as it gives valuable insights beyond the specific application of our work.
@@
  • Finally, in Section 4, our experimental results further validate the different performance aspects of Neural-ADB, highlighting its sample efficiency for preference data collection.
@@ 2 Problem Setting
We model active human preference feedback collection as an active contextual dueling bandit problem, where a labeler (human or simulator) provides preference feedback for a chosen pair of arms.
Active contextual dueling bandit.   We consider an active contextual dueling bandit problem, where the underlying latent reward function can be non-linear. In each iteration of this problem, the learner’s goal is to select a triplet containing a context and two arms for collecting preference feedback from a labeler/human such that the collected preference dataset leads to superior performance. Let $\mathcal{C}$ be the set of contexts and $\mathcal{A}$ be the set of all possible arms. In each iteration, the learner selects a context $c_{t}\in\mathcal{X}$ and then two arms (denoted as $a_{t,1}$ and $a_{t,2}$) from the set of arms $\mathcal{A}$. After selecting the triplet of context and two arms, the learner receives a stochastic preference feedback $y_{t}$, where $y_{t}=1$ implies the arm $a_{t,1}$ is preferred over arm $a_{t,2}$ for the context $c_{t}$ and $y_{t}=0$ otherwise. We use $\varphi(c_{t},a)$ to denote the context-arm feature vector for context $c_{t}$ and an arm $a$, where $\varphi:\mathcal{C}\times\mathcal{A}\rightarrow\mathbb{R}^{d}$ is a known feature map, such as one that concatenates the context and arm features.
@@
  • Let $\kappa_{\mu}\doteq\inf\limits_{c\in\mathcal{C},a,b\in\mathcal{A}}\dot{\mu}(f(\varphi(c,a))-f(\varphi(c,b)))>0$ for all triplets of context $(c)$ and pair of arms $(a,b)$.
@@
Performance measure.  We denote the collected preference dataset up to $T$ iterations by $\mathcal{D}_{T}=\{(c_{s},a_{s,w},a_{s,l},y_{s})\}_{s=1}^{T}$, where $a_{s,w}\succ a_{s,l}$ for the selected context $c_{s}$ in iteration $s$. We aim to learn a policy, $\pi:\mathcal{C}\rightarrow\mathcal{A}$ from the collected preference dataset $\mathcal{D}_{T}$ that achieves the worst sub-optimality gap across all contexts in $\mathcal{C}$, which is defined as follows:
@@ 3 Algorithm for Active Human Preference Feedback Collection
In this section, we introduce Neural-ADB, a simple yet principled and practical algorithm designed to efficiently select context-arm pairs for collecting preference feedback. Neural-ADB consists of two main components: context selection and arm selection. Since the arm selection strategy depends on the estimated reward function, we first explain how an NN can be used to estimate the unknown reward function. We will then give details of the context and arm selection strategies, followed by our theoretical results that validate the effectiveness of Neural-ADB.
@@ 3.1 Reward function estimation using neural network
where $\text{ReLU}(v)=\max\{v,0\}$, $\bm{W}_{1}\in\mathbb{R}^{w\times d}$, $\bm{W}_{l}\in\mathbb{R}^{w\times w}$ for $2\leq l<D$, $\bm{W}_{D}\in\mathbb{R}^{w\times 1}$. The parameters of the NN are represented by $\theta=\left(\text{vec}\left(\bm{W}_{1}\right);\cdots\text{vec}\left(\bm{W}_{D}\right)\right)$, where $\text{vec}\left(A\right)$ transforms an $m\times n$ matrix $A$ into a vector of dimension $mn$. We use $p$ to represent the total number of NN parameters, which is given by $p=dw+w^{2}(D-1)+w$, and $g(x;\theta)$ to denote the gradient of NN $h(x;\theta)$ with respect to $\theta$. At the end of each iteration $t$, the preference dataset $\mathcal{D}_{t}=\left\{(c_{s},a_{s,w},a_{s,l},y_{s})\right\}_{s=1}^{t}$ is used to estimate the reward function $f$ by training an NN $h$ (parameterized by $\theta_{t+1}$) using gradient descent to minimize the following binary cross entropy loss function:
@@
where $\theta_{0}$ denotes the initial parameter of the NN that is initialized according to the standard practice in neural bandits (Zhou et al., 2020; Zhang et al., 2021) (see Algorithm 1 in Zhang et al. (2021) for details). Minimizing the first term in the above loss function (that involves the summation over the $t$ terms) corresponds to finding the maximum log-likelihood estimate of the parameters $\theta$.
@@ 3.2 Neural-ADB
We next propose a simple yet principled and practical algorithm, Neural-ADB, that consists of two key components: Context selection and arm selection. Neural-ADB works as follows: At the beginning of the iteration $t$, we first select the context as follows:
@@
where $V_{t-1}=\frac{\lambda}{\kappa_{\mu}}\mathbb{I}_{p}+\sum_{s=1}^{t-1}z_{s}{z_{s}}^{\top}\frac{1}{w}$ in which $z_{s}=\varphi(c_{s},a_{s,w})-\varphi(c_{s},a_{s,l})=g(\varphi(c_{s},a_{s,w});\theta_{0})-g(\varphi(c_{s},a_{s,l});\theta_{0})$, and $g(\varphi(c_{s},a_{s,i});\theta_{0})/\sqrt{w}$ is used as the Random features approximation for the context-arm feature vector $\varphi(c_{s},a_{s,i})$. This strategy is adapted from the context selection strategy [^3] Note that selecting contexts uniformly at random suffer a constant sub-optimality gap (Das et al., 2024, Theorem 3.2). from Das et al. (2024). After selecting context $c_{t}$, Neural-ADB uses the trained NN (as an estimate of the unknown reward function) to decide which two arms must be selected. To do so, Neural-ADB uses UCB- and TS-based arm selection strategies, which efficiently balance the trade-off between exploration and exploitation (Lattimore and Szepesvári, 2020) due to the bandit nature of preference feedback, as preference feedback is only observed for the selected pair of arms.
@@
The second arm $a_{t,2}$ is selected optimistically by maximizing the UCB value as follows:
@@
$\sigma_{t-1}^{2}(c,a,b)\doteq\frac{\lambda}{\kappa_{\mu}}\left\|\frac{1}{\sqrt{w}}(\varphi(c,a)-\varphi(c,b))\right\|^{2}_{V_{t-1}^{-1}}.$ (6)
A larger value of $\sigma_{t-1}^{2}(c_{t},a_{t,1},b)$ implies that arm $b$ is significantly different from $a_{t,1}$, given the contexts and arm pairs already selected. As a result, the second term in Eq. 5 makes the second arm different from the first arm which ensures exploration.
TS-based arm selection strategy.   Thompson sampling (TS) selects an arm based on its probability of being the best (Thompson, 1933). Several works (Chapelle and Li, 2011; Agrawal and Goyal, 2013; Chowdhury and Gopalan, 2017; Li et al., 2024) have shown that TS empirically outperforms UCB-based bandit algorithms. Therefore, we also propose a TS-based arm selection strategy in which the first arm is also selected using Eq. 4 and the second arm is selected differently. To select the second arm, it first samples a score $s_{t}(b)\sim\mathcal{N}\big{(}h(\varphi(c_{t},b);\theta_{t})-h(\varphi(c_{t},a_{t,1});\theta_{t}),\nu_{T}^{2}\sigma_{t-1}^{2}(c_{t},a_{t,1},b)\big{)}$ for every arm $b\in\mathcal{A}\setminus\{a_{t,1}\}$ and then selects the second arm that maximizes the samples scores as follows:
@@ 3.3 Theoretical Results
Equipped with 1, 1, and 2, we will now provide an upper bound on the worse sub-optimality gap for a policy learned by Neural-ADB while using UCB- and TS-based arm selection strategy for a given context.
@@
The proof follows by applying 2, setting $z=\varphi(c,a^{\star})-\varphi(c,\pi(c))$ in Eq. 6, and then using 1. Note that $\varepsilon^{\prime}_{w,T}=O(1/T)$ and $\widetilde{d}=\widetilde{o}(\sqrt{T})$ as long as the NN width $w$ is large enough (Zhou et al., 2020; Zhang et al., 2021; Verma et al., 2025). Above 2 and 3 show that the worst sub-optimality gap of the policy learned by Neural-ADB with UCB- and TS-based arm selection strategies decreases at a sub-linear rate with respect to the size of preference dataset, specifically at rate of $\tilde{O}((\tilde{d}/T)^{\frac{1}{2}})$, where $\tilde{O}$ hides the logarithmic factors and constants. Further, the decay rate of the worst sub-optimality gap for Neural-ADB improves by a factor of $\tilde{O}(({\tilde{d}\log T})^{\frac{1}{2}})$ compared to exiting algorithms (Mehta et al., 2023; Das et al., 2024), thereby bridging the gap between theory and practice.